OpenStudy (anonymous):

Solve the equation: 4x^2 = 50

OpenStudy (anonymous):

4x^2=50 x^2=50/4 x^2=12.5 \[x=\sqrt{12.5}\] x= 3.5

OpenStudy (anonymous):

it says make sure to simplify radicals, is this still the right answer?

OpenStudy (anonymous):

x = sqrt ( 12.5) = sqrt ( 25 * .5) = 5 sqrt (.5)